Dome Metals has credit sales of $180,000yearly with credit terms of net 60 days, which is also the average collection period. Dome does not offer a discount for early payment, so its customers take the full 60 dys to pay. What is the average receivables balance? Receivables turnover?
Answer #1

Part A:

Average Daily Sales = Sales/360 Days

Average Daily Sales = 180000/360 = $500

Average Accounts Receivables Balance = Average Daily Sales*60 = 500*60 = $30000

Answer for Part A is $30000.

----------

Part B:

Receivables Turnover = Credit Sales/Average Receivables Balance = 180000/30000 = 6

Answer for Part B is 6.

Thanks.